OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

y = 2x + 6 y = 1/2x + 3 y = 1/2x + 6 y = 2x + 3